# ResiDual Transformer Alignment with Spectral Decomposition

### v1

@@ Abstract
When examined through the lens of their residual streams, a puzzling property emerges in transformer networks: residual contributions (e.g., attention heads) sometimes specialize in specific tasks or input attributes. In this paper, we analyze this phenomenon in vision transformers, focusing on the spectral geometry of residuals, and explore its implications for modality alignment in vision-language models. First, we link it to the intrinsically low-dimensional structure of visual head representations, zooming into their principal components and showing that they encode specialized roles across a wide variety of input data distributions. Then, we analyze the effect of head specialization in multimodal models, focusing on how improved alignment between text and specialized heads impacts zero-shot classification performance. This specialization-performance link consistently holds across diverse pre-training data, network sizes, and objectives, demonstrating a powerful new mechanism for boosting zero-shot classification through targeted alignment. Ultimately, we translate these insights into actionable terms by introducing ResiDual, a technique for spectral alignment of the residual stream. Much like panning for gold, it lets the noise from irrelevant unit principal components (i.e., attributes) wash away to amplify task-relevant ones. Remarkably, this dual perspective on modality alignment yields fine-tuning level performances on different data distributions while modeling an extremely interpretable and parameter-efficient transformation, as we extensively show on more than 50 (pre-trained network, dataset) pairs.
00footnotetext: Correspondence to: lorenzo.basile@phd.units.it and maiorca@di.uniroma1.it
@@ 1 Introduction
Figure 1: From the transformer’s residual stream, we can obtain direct head contributions from across the network. In a multimodal, zero-shot classification setting (e.g., in CLIP) task boundaries originate from text that can be at different level of concept granularity. When some heads are specialized in some properties (e.g., shape, pattern, color), they can be better aligned to apply some boundaries than the original output. In this example, only the fine-grained task (brown) can correctly separate the samples at the output level.
@@
In modern transformer networks, the model’s final output is produced by applying a simple linear transformation to the residual stream (up to LayerNorm). This residual stream accumulates information additively, drawing from each attention head and all MLP layers (Elhage et al., 2021), producing a general-purpose representation used as a feature set for many tasks. This decomposition raises an intriguing question: are all these units essential for solving specific tasks, or do some introduce noise that obscures task-relevant information? Typically, there’s a trade-off between a model’s generalization and its performance on specific tasks. However, it might be that to specialize a model for a specific task, we don’t need to retrain the whole model. Instead, by manipulating the residual stream, we can boost the units already aligned with the task, amplifying relevant signals while reducing noise.
@@
In this paper, we tackle this question from the perspective of the latent geometry of residual units. First, on a variety of transformer-based vision models, including multiple versions of CLIP (Radford et al., 2021), BLIP (Li et al., 2022), ViT (Dosovitskiy et al., 2021) and DINOv2 (Oquab et al., 2024), we show that such units are embedded in low-dimensional manifolds and that, when there is specialization, it can be traced back to the role of few principal components. Then, by introducing a spectral analysis method based on a discrete version of principal angles (Björck & Golub, 1973), we quantitatively measure the similarity of residual units across different datasets revealing that the roles of specialized units remain surprisingly stable.
@@
Building on this insight, we hypothesize that, in many cases, the information necessary for solving a task is already embedded within a subset of highly specialized residual units. We show that this picture emerges clearly in vision-language models like CLIP, where we find units or sets of units that align with textual attributes more precisely than the full model output, on a given task. In fact, as the output combines all residual units, this relevant information may be obfuscated by other units that introduce irrelevant signals. Instead of fine-tuning the entire model, we propose to isolate and enhance the task-relevant units by filtering out the noise - akin to panning for gold. By doing so, we can significantly boost model performance with up to 4 orders or magnitude less parameters than full fine-tuning and 2 less than those needed for training a simple linear transformation at the output level. To implement this, we introduce ResiDual, a novel approach that focuses on the principal components (PCs) of the residual units to identify and retain the needed information. This framework selectively reweights the most relevant PCs, amplifying the signals that align with the task objective while remaining computationally efficient. This spectral reweighting of individual units addresses nonlinear interactions between them, and provides a geometrically principled and interpretable method for optimizing transformer models by capitalizing on the knowledge they already possess.
@@
  • We identify task-specific units in vision-language models, showcasing that focusing on these units in zero-shot settings can outperform using the full residual output when there is latent alignment between units and tasks.
  • We present ResiDual, a geometrically grounded method for manipulating transformers by reweighting the most relevant principal components of the residual units. This approach can sidestep the need for full-model finetuning, as it reaches competitive performance with minimal parameter overhead.
@@ 2 Related Work
Transformer networks (Vaswani et al., 2017) rely on residual connections around each multi-head attention (MHA) and MLP layer, resulting in a final representation that combines contributions from all units across layers by simple summation. Techniques like logit lens (nostalgebraist, 2020), Direct Logit Attribution (DLA) (Elhage et al., 2021) – and Cancedda (2024), at the spectral level – focus on how individual layers or residual units (such as MLPs or attention heads) affect the final output in logit space, given that their contributions are projected upstream via linear transformations (up to LayerNorm (Lei Ba et al., 2016), an affine one). Here, we apply this residual decomposition to provide a more comprehensive understanding of how these units interact and align across different tasks, revealing the deeper structure within the residual space.
@@
To understand the geometric nature of latent manifolds, previous works analyze the intrinsic dimensionality ($I_{d}$) (Ansuini et al., 2019; Cheng et al., 2023; Valeriani et al., 2024) of the representations within the network, which is typically much lower than the embedding dimension. We posit that the Linear Representation Hypothesis (Park et al., ; Jiang et al., 2024), which suggests that transformer representations encode high-level concepts in linear directions, complements this view hinting at transformer models sometimes modulating input attributes in a linearly structured and low-dimensional (i.e. specialized) way. Previous works in language modeling have highlighted this specialization (Voita et al., 2019; Michel et al., 2019; Li et al., 2023; Lv et al., 2024; Chughtai et al., 2024), revealing that only a few attention heads are responsible for specific tasks and that they assume specialized and interpretable roles. Our analysis bridges geometry and specialization, revealing that vision transformer heads are low-dimensional (though often nonlinear, especially in deeper layers) and highly specialized for downstream tasks.
@@ 3 The Geometry of Residual Units
${\bm{Y}}=\sum_{i=1}^{|{\mathbb{U}}|}{\bm{U}}_{i}={\bm{X}}_{0}+\sum_{i=1}^{|{\mathbb{H}}|}{\bm{H}}_{i}+\sum_{j=1}^{|{\mathbb{M}}|}{\bm{M}}_{j}$ (1)
With ${\bm{Y}}$ as the final output of the model, summing up all the residual units in ${\bm{U}}\in{\mathbb{U}}$. Please refer to Section A.1 for a more rigorous description of the residual decomposition.
@@ 3.1 Residual Dimensionality
Despite being embedded into a higher-dimensional space, head representations exhibit an even lower intrinsic dimensionality ($I_{d}$) than that of the original ambient space. This indicates that, irrespective of their high-dimensional embedding, the essential structure of head representations is highly compressed and governed by a compact, low-dimensional geometry. In short, the intrinsic dimensionality of a dataset is the least number of variables required to satisfactorily describe the data points. Ideally, if data lie on a linear manifold (a hyperplane), the intrinsic dimensionality coincides with the number of principal components required to completely explain their variance. In a more realistic setting, data lie on curved, nonlinear manifolds, and linear estimators like PCA fail to capture their real intrinsic dimensionality. In such cases, one can resort to nonlinear $I_{d}$ estimators. Among them, we choose to employ the TwoNN (Facco et al., 2017) because of its efficiency and stability on complex and non-uniform manifolds.
@@
We start by evaluating the intrinsic dimensionality of head representations across multiple transformer-based vision architectures pre-trained with different objectives (supervised, unsupervised, self-supervised). Namely, we employ OpenAI’s CLIP (Radford et al., 2021), OpenCLIP (Cherti et al., 2023), BLIP (Li et al., 2022), ViT (Dosovitskiy et al., 2021) and DINOv2 (Oquab et al., 2024), all in their version based on ViT-Large (results on ViT-Base models are in the Appendix in Figure 7). We feed them a subset of the training set of ImageNet (Russakovsky et al., 2015) containing 80000 images stratified on the class labels and we extract the representations for all attention heads. Then, we compute the intrinsic dimensionality of such representations using a linear estimator (PCA) and a nonlinear one (TwoNN). Linear $I_{d}$ is computed as the number of components needed by PCA to explain 99% of head variance.
Figure 2: Heads in early layers show low-dimensional, linear structures, as suggested by similar intrinsic dimension estimates from PCA (L) and TwoNN (N). Moving toward the output layer, the nonlinear dimensionality peaks and then decreases, while PCA’s linear estimate continues to rise, indicating increasing nonlinearity in head manifolds (Ratio = $\frac{L}{N}$). The first principal component (EVR1) explains around 50% of the variance in early layers, dropping to around 10% in later layers.
@@
We report our results in Figure 2. We observe that the true head dimensionality (the one computed with a nonlinear estimator, TwoNN) tends to increase in the first half of the model and to decrease towards the last few layers, following a characteristic hunchback shape, similar to previous findings in other vision architectures (Ansuini et al., 2019). However, the number of dimensions returned by the linear estimator grows constantly through the model. This disparity, witnessed by the increasing ratio between the two estimates, indicates that the units of the first layers are close to linear, while in late layers they lie on more curved manifolds. The last column shows the average explained variance ratio (EVR) of the first PCA component and highlights that heads in the first layers are largely explained by this direction, while it still accounts for a nontrivial 10% of head variance in late layers.
@@ 3.2 Principal Components Encode Unit Semantics
The low dimensionality of head encodings results in relevant consequences for their interpretability. Head representations can be easily approximated with sparse recovery algorithms, in a way that is akin of performing PCA, but over a discrete set of vectors. For CLIP models, this approach has been recently explored by Gandelsman et al. (2024). There, the authors introduce a sparse approximation algorithm, TextSpan (TS), and decompose head encodings using a set of textual descriptions, coming from the text branch of CLIP, as a dictionary. They observe strong specialization properties, witnessed by high coherence in the textual explanations of each head. We link TextSpan to the more established family of Matching Pursuit (MP) (Mallat & Zhang, 1993) algorithms, widely employed in signal processing. More specifically, as we show in the Section A.2, TextSpan is analogous to Simultaneous Orthogonal Matching Pursuit (SOMP) (Tropp et al., 2006), with light modifications. TextSpan, like any MP algorithm, approximates the signal through linear combinations of basis functions. Considering the high nonlinearity of later layers (Section 3.1), and TS being a linear sparse approximation method, we now want to investigate whether TS is, in reality, focusing on the first principal components of the signal (head-level representations).
@@
For this experiment, we position ourselves in the same setup of the original TS paper (Gandelsman et al., 2024). Hence, we consider the attention heads belonging to the last 4 layers of OpenCLIP-L. We use two sparse approximation algorithms: the original TextSpan, which operates on the whole head representation, and Orthogonal Matching Pursuit (OMP) (Pati et al., 1993). Different from TextSpan, OMP computes sparse approximations of vectors, not matrices (like head representations). Therefore, in this experiment, we apply OMP to the first principal component of each head. We denote this method as OMP1. The dictionary we use contains the encodings produced by OpenCLIP-L for the set of image descriptions provided by Gandelsman et al. (2024). We apply both algorithms to select 5 descriptions for each head and compute an agreement score between the two sets. The agreement is computed as the absolute Z-score of the cosine similarity (sim) between them, compared with the average cosine similarity $\mu$ between the descriptions selected by TextSpan and the entire dictionary $Z=\frac{|\overline{\texttt{sim}(\text{TS},\text{OMP}_{1})}-\mu|}{\sigma}$.
@@
We report in the left panel of Figure 3 the agreement scores. The right panel reports a few examples (one per layer) of descriptions obtained using the two algorithms. We observe that a high Z-score (e.g., head 8 of layer 22, which is almost $5\sigma$ away from $\mu$), is reflected in extremely similar descriptions from the two methods. When the agreement is lower, as in the case of head 20 from layer 8, even though they share some high-level semantics, the two sets of descriptions substantially differ.
@@
Figure 3: Comparison between TextSpan and Ortogonal Matching Pursuit on the first principal component (OMP1), applied to the heads of OpenCLIP-L. Left: agreement score between the descriptions returned by the two methods. Right: qualitative comparison of selected descriptions for 4 heads, one per layer, at different agreement levels. A similar analysis for the second principal component is presented in the Appendix in Figure 8.

### v2

@@ Abstract
When examined through the lens of their residual streams, a puzzling property emerges in transformer networks: residual contributions (e.g., attention heads) sometimes specialize in specific tasks or input attributes. In this paper, we analyze this phenomenon in vision transformers, focusing on the spectral geometry of residuals, and explore its implications for modality alignment in vision-language models. First, we link it to the intrinsically low-dimensional structure of visual head representations, zooming into their principal components and showing that they encode specialized roles across a wide variety of input data distributions. Then, we analyze the effect of head specialization in multimodal models, focusing on how improved alignment between text and specialized heads impacts zero-shot classification performance. This specialization-performance link consistently holds across diverse pre-training data, network sizes, and objectives, demonstrating a powerful new mechanism for boosting zero-shot classification through targeted alignment. Ultimately, we translate these insights into actionable terms by introducing ResiDual, a technique for spectral alignment of the residual stream. Much like panning for gold, it lets the noise from irrelevant unit principal components (i.e., attributes) wash away to amplify task-relevant ones. Remarkably, this dual perspective on modality alignment yields fine-tuning level performance on different data distributions while modelling an extremely interpretable and parameter-efficient transformation, as we extensively show on 70 pre-trained network-dataset combinations (7 models, 10 datasets).
00footnotetext: Equal contribution. Work done while visiting ISTA. Code is available at https://github.com/Flegyas/ResiDual
@@ 1 Introduction
Figure 1: From the transformer’s residual stream, direct contributions from individual heads across the network can be analyzed. In a multimodal, zero-shot classification setting (e.g., in CLIP), task boundaries are defined by text prompts that may vary in their conceptual granularity. When certain heads are specialized in particular features (e.g., shape, pattern, color), they may more accurately apply these boundaries than the model’s original output. In this example, only the fine-grained task (brown) effectively separates the samples at the output level.
@@
In modern transformer networks, the model’s final output is produced by applying a simple linear transformation to the residual stream (up to LayerNorm). This residual stream accumulates information additively, drawing from each attention head and all MLP layers (Elhage et al., 2021), producing a general-purpose representation used as a feature set for many tasks. This decomposition raises an intriguing question: are all these units essential for solving specific tasks, or do some introduce noise that obscures task-relevant information? Typically, there is a trade-off between a model’s generalization and its performance on specific tasks. However, it might be that to specialize a model for a specific task, we do not need to retrain the whole model. This is how we can define task-irrelevant and noise units: task-irrelevant units are those that can be deactivated without affecting performance, while noise refers to units that, when deactivated, enhance performance by removing signals misaligned with the task. By manipulating the residual stream, we can boost the units already aligned with the task, amplifying relevant signals while reducing noise.
Contribution
@@
In this paper, we tackle this question from the perspective of the latent geometry of residual units. First, on a variety of transformer-based vision models, including multiple versions of CLIP (Radford et al., 2021), BLIP (Li et al., 2022), ViT (Dosovitskiy et al., 2021) and DINOv2 (Oquab et al., 2024), we show that such units are embedded in low-dimensional manifolds and that, when there is specialization, it can be traced back to the role of few principal components. Then, by introducing a spectral analysis method based on a discrete version of principal angles (Björck & Golub, 1973), we quantitatively measure the similarity of residual units across different datasets, revealing that the roles of specialized units remain surprisingly stable.
@@
Building on this insight, we hypothesize that, in many cases, the information necessary for solving a task is already embedded within a subset of highly specialized residual units. We show that this picture emerges clearly in vision-language models like CLIP, where we find units or sets of units that align with textual attributes more precisely than the full model output on a given task. In fact, as the output combines all residual units, this relevant information may be obfuscated by other units that introduce irrelevant signals. Instead of fine-tuning the entire model, we propose to isolate and enhance the task-relevant units by filtering out the noise – akin to panning for gold. By doing so, we can significantly boost model performance with up to 4 orders of magnitude fewer parameters than full fine-tuning and 2 less than those needed for training a simple linear transformation at the output level.
To implement this, we introduce ResiDual, a novel approach that focuses on the principal components (PCs) of the residual units to identify and retain the needed information. This framework selectively reweights the most relevant PCs, amplifying the signals that align with the task objective while remaining computationally efficient. This spectral reweighting of individual units addresses nonlinear interactions between them and provides a geometrically principled and interpretable method for optimizing transformer models by capitalizing on the knowledge they already possess.
@@
  • We identify task-specific units in vision-language models, showcasing that focusing on these units in zero-shot settings can outperform using the full residual output when there is latent alignment between units and tasks;
  • We introduce ResiDual, a geometrically grounded method for manipulating transformers by reweighting the most relevant principal components of the residual units. This approach can sidestep the need for full-model finetuning, as it reaches competitive performance with minimal parameter overhead.
@@ 2 Related Work
Transformer networks (Vaswani et al., 2017) rely on residual connections around each multi-head attention (MHA) and MLP layer, resulting in a final representation that combines contributions from all units across layers by simple summation. Techniques like logit lens (nostalgebraist, 2020) and Direct Logit Attribution (DLA) (Elhage et al., 2021) – and Cancedda (2024), at the spectral level – focus on how individual layers or residual units (such as MLPs or attention heads) affect the final output in logit space, given that their contributions are projected upstream via linear transformations (up to LayerNorm (Lei Ba et al., 2016), an affine one). Here, we apply this residual decomposition to provide a more comprehensive understanding of how these units interact and align across different tasks, revealing the deeper structure within the residual space.
@@
To understand the geometric nature of latent manifolds, previous works analyze the intrinsic dimensionality ($I_{d}$) (Ansuini et al., 2019; Cheng et al., 2023; Valeriani et al., 2024) of the representations within the network, which is typically much lower than the embedding dimension. We posit that the Linear Representation Hypothesis (Park et al., ; Jiang et al., 2024), which suggests that transformer representations encode high-level concepts in linear directions, complements this view hinting at transformer models sometimes modulating input attributes in a linearly structured and low-dimensional (i.e., specialized) way. Previous works in language modeling have highlighted this specialization (Voita et al., 2019; Michel et al., 2019; Li et al., 2023; Lv et al., 2024; Chughtai et al., 2024), revealing that only a few attention heads are responsible for specific tasks and that they assume specialized and interpretable roles. Our analysis bridges geometry and specialization, revealing that vision transformer heads are low-dimensional (though often nonlinear, especially in deeper layers) and highly specialized for downstream tasks.
@@ 3 The Geometry of Residual Units
${\bm{Y}}=\sum_{i=1}^{|{\mathbb{U}}|}{\bm{U}}_{i}={\bm{X}}_{0}+\sum_{i=1}^{|{\mathbb{H}}|}{\bm{H}}_{i}+\sum_{j=1}^{|{\mathbb{M}}|}{\bm{M}}_{j}\,,$ (1)
with ${\bm{Y}}$ as the final output of the model, summing up all the residual units in ${\bm{U}}\in{\mathbb{U}}$. Please refer to Section A.1 for a more rigorous description of the residual decomposition. In this paper, we focus on the geometry of attention head representations, as prior works (e.g., Voita et al. (2019); Gandelsman et al. (2024)) have demonstrated their specialization in distinct tasks. MLP layers mix outputs from multiple heads by design, leading to more entangled and less interpretable representations (as we show in Figure 11).
@@ 3.1 Residual Dimensionality
Despite being embedded into a higher-dimensional space, head representations exhibit an even lower intrinsic dimensionality ($I_{d}$) than that of the original ambient space. This indicates that irrespective of their high-dimensional embedding, the essential structure of head representations is highly compressed and governed by a compact, low-dimensional geometry. In short, the intrinsic dimensionality of a dataset is the least number of variables required to satisfactorily describe the data points. Ideally, if data lie on a linear manifold (a hyperplane), the intrinsic dimensionality coincides with the number of principal components required to completely explain their variance. In a more realistic setting, data lie on curved, nonlinear manifolds, and linear estimators like PCA fail to capture their real intrinsic dimensionality. In such cases, one can resort to nonlinear $I_{d}$ estimators. Among them, we choose to employ the TwoNN (Facco et al., 2017) because of its efficiency and stability on complex and non-uniform manifolds.
@@
We start by evaluating the intrinsic dimensionality of head representations across multiple transformer-based vision architectures pre-trained with different objectives (supervised, unsupervised, self-supervised). Namely, we employ OpenAI’s CLIP (Radford et al., 2021), OpenCLIP (Cherti et al., 2023), BLIP (Li et al., 2022), ViT (Dosovitskiy et al., 2021) and DINOv2 (Oquab et al., 2024), all in their version based on ViT-Large (results on ViT-Base models are in the Appendix in Figure 8). We feed them a subset of the training set of ImageNet (Russakovsky et al., 2015) containing 80000 images stratified on the class labels, and we extract the representations for all attention heads. Then, we compute the intrinsic dimensionality of such representations using a linear estimator (PCA) and a nonlinear one (TwoNN). Linear $I_{d}$ is computed as the number of components needed by PCA to explain 99% of head variance.
Figure 2: Heads in early layers show low-dimensional, linear structures, as suggested by similar intrinsic dimension estimates from PCA (L) and TwoNN (N). Moving toward the output layer, the nonlinear dimensionality peaks and then decreases, while PCA’s linear estimate continues to rise, indicating increasing nonlinearity in head manifolds ($\text{Ratio}=\frac{L}{N}$). The first principal component (EVR1) explains around 50% of the variance in early layers, dropping to around 10% in later layers.
@@
We report in Figure 2 the results for head units on ImageNet (a subset of 80000 images stratified by class), in Figure 10 the results for other datasets and in Figure 11 those for MLP units on ImageNet. We observe that the true head dimensionality (the one computed with a nonlinear estimator, TwoNN) tends to increase in the first half of the model and to decrease towards the last few layers, following a characteristic hunchback shape, similar to previous findings in other vision architectures (Ansuini et al., 2019). However, the number of dimensions returned by the linear estimator grows constantly through the model. This disparity, reflected in the growing ratio between the two estimates, suggests that the units in the early layers are close to linear, while those in the later layers lie on more curved manifolds. The last column shows the average explained variance ratio (EVR) of the first PCA component and highlights that heads in the first layers are largely explained by this direction, while it still accounts for a nontrivial 10% of head variance in late layers.
@@ 3.2 Principal Components Encode Unit Semantics
The low dimensionality of head encodings results in relevant consequences for their interpretability. Head representations can be easily approximated with sparse recovery algorithms in a way that is akin to performing PCA, but over a discrete set of vectors. For CLIP models, this approach has been recently explored by Gandelsman et al. (2024). There, the authors introduce a sparse approximation algorithm, TextSpan (TS), and decompose head encodings using a set of textual descriptions coming from the text branch of CLIP as a dictionary. They observe strong specialization properties, witnessed by high coherence in the textual explanations of each head.
We link TextSpan to the more established family of Matching Pursuit (MP) (Mallat & Zhang, 1993) algorithms, widely employed in signal processing. More specifically, as we show in the Section A.2, TextSpan is analogous to Simultaneous Orthogonal Matching Pursuit (SOMP) (Tropp et al., 2006), with light modifications. TextSpan, like any MP algorithm, approximates the signal through linear combinations of basis functions. Considering the high nonlinearity of later layers (Section 3.1), and TS being a linear sparse approximation method, we now want to investigate whether TS is, in reality, focusing on the first principal components of the signal (head-level representations).
@@
For this experiment, we position ourselves in the same setup as the original TS paper (Gandelsman et al., 2024). Hence, we consider the attention heads belonging to the last 4 layers of OpenCLIP-L. We use two sparse approximation algorithms: the original TextSpan, which operates on the whole head representation, and Orthogonal Matching Pursuit (OMP) (Pati et al., 1993). Different from TextSpan, OMP computes sparse approximations of vectors, not matrices (like head representations). Therefore, in this experiment, we apply OMP to the first principal component of each head. We denote this method as OMP1. The dictionary we use contains the encodings produced by OpenCLIP-L for the set of image descriptions provided by Gandelsman et al. (2024). We apply both algorithms to select 5 descriptions for each head and compute an agreement score between the two sets. The agreement is computed as the absolute Z-score of the cosine similarity (sim) between them, compared with the average cosine similarity $\mu$ between the descriptions selected by TextSpan and the entire dictionary $Z=\frac{|{\texttt{sim}(\text{TS},\text{OMP}_{1})}-\mu|}{\sigma}$.
@@
We report in the left panel of Figure 3 the agreement scores. The right panel reports a few examples (one per layer) of descriptions obtained using the two algorithms. We observe that a high Z-score (e.g., head 8 of layer 22, which is almost $5\sigma$ away from $\mu$), is reflected in extremely similar descriptions from the two methods. When the agreement is lower, as in the case of head 20 from layer 8, the two sets of descriptions substantially differ even though they share some high-level semantics.
@@
Figure 3: Comparison between TextSpan and Orthogonal Matching Pursuit on the first principal component (OMP1), applied to the heads of OpenCLIP-L. Left: agreement score between the descriptions returned by the two methods. Right: qualitative comparison of selected descriptions for 4 heads, one per layer, at different agreement levels. A similar analysis for the second principal component is presented in the Appendix in Figure 9.